The next 3 questions involve the following supply and demand equations (please use diagrams to explain)

Supply:q= (1/4)p

Demand:q= 6−(1/2)p

17.The government enacts a price floor of $10. What is the surplus?

(A) 2.5.

(B) 1.5.

(C) 3.5.

(D) None of the above

What is the Deadweight Loss under a price floor of $10?

(A) $2.

(B) $3.

(C) $6.

(D) None of the above.
Answer #1

Ans) 1) At price of $10÷

Qd= 6 - (1/2)P = 6 - (1/2)×10 = 1

Qs= (1/4)P = (1/4)×10 = 2.5

Surplus = Qs - Qd = 2.5 - 1 = 1.5

Option b.

2) Before finding deadweightloss, lets draw the graph,

At Q= 0

Qs= 1/4P

0=1/4P

P=0.

And again at Q=0

Qd= 6 - 1/2P

0= 6 - 1/2 P

P= 12

Here we have got one coordinate for supply curve i.e (0,0)

And one for demand curve i.e (0,12)

Now, at equilibrium Qd=Qs

6-(1/2)P = (1/4)P

P= 8

To determine equilibrium quantity, put value of P in Qd or Qs

Qd= 6-(1/2)P = 6-(1/2)8 = 2

Qs= (1/4)P = (1/4)8 = 2

So, here we get another coordinate for demand curve and supply curve i.e (2,8)

لا و مص ما ک له 15 2 2.5 3 55

Deadweightloss = 1/2× base × height = 1/2× (2 -1)×(10-4) = 3

Option b.